OpenStudy (anonymous):

The number of ways the letters of the alphabet can be placed in a straight line with the restriction that the letters of the sentence UP THE BIG SKY WORLD must not be together is: A 26! − 11!16! B 26! + 16! C 16! − 8! D 6!16! E 6!16!26!

ganeshie8 (ganeshie8):

how many letters are there in the sentence

ganeshie8 (ganeshie8):

UP THE BIG SKY WORLD

ganeshie8 (ganeshie8):

they are 16 and all of them are uniq

ganeshie8 (ganeshie8):

26 letters can be arranged in 26!

ganeshie8 (ganeshie8):

out of these permutations (26-16+1) x 16! will have the letters in given sentence together

ganeshie8 (ganeshie8):

26! - (26-16+1) x 16! permutations will not have the letters in given sentence all together

OpenStudy (anonymous):

can you explain your second last point?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

16 letters are to be all together

ganeshie8 (ganeshie8):

26 letters are divided into one group of 16 letters + rest of the letters

OpenStudy (anonymous):

yep k

ganeshie8 (ganeshie8):

total 17 groups

OpenStudy (anonymous):

kk

ganeshie8 (ganeshie8):

ok forget it

OpenStudy (anonymous):

ok, i got it now.

ganeshie8 (ganeshie8):

actually its not 17 groups, 1 group and (26-11) letters = 1 + 10 = 11 groups

OpenStudy (anonymous):

yeh i thought that was weird and above u did correctly.

ganeshie8 (ganeshie8):

11 groups will form 11! permutations each of the permutation will have 16! permutations for the given sentence

ganeshie8 (ganeshie8):

yeah 17! is misleading. i messed up there. excuse that :D